OpenStudy (anonymous):

Solve the problem. The formula A = P(1 + r)^2 is used to find the amount of money, A, in an account after P dollars have been invested in the account paying an annual interest rate, r, for 2 years. Find the interest rate r if $500 grows to $980 in 2 years. Seleccione una respuesta. a. 96% b. 4% c. 240% d. 40%

OpenStudy (anonymous):

980=500(1+r)^2 divide by 500 to get 1.96=(1+r)^2 take the square root of both sides to get 1.4=1+r. Subtract 1 to get r=.4 and that in percent is 40% so answer choice D is correct.

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

Yes, if you aren’t just plug it back in to the equation and see if I’m right.

OpenStudy (anonymous):

40 % is correct; i made a mistake earlier; my apologies